# Many-Tier Instruction Hierarchy in LLM Agents

### 10 Apr 2026 v1

@@ Abstract
Large language model agents receive instructions from many sources—system messages, user prompts, tool outputs, and more—each carrying different levels of trust and authority. When these instructions conflict, models must reliably follow the highest-privilege instruction to remain safe and effective. The dominant paradigm, instruction hierarchy (IH), assumes a fixed, small set of privilege levels (typically fewer than five) defined by rigid role labels (e.g., system > user). This is inadequate for real-world agentic settings, where conflicts can arise across far more sources and contexts. In this work, we propose Many-Tier Instruction Hierarchy (ManyIH), a paradigm for resolving instruction conflicts among instructions with arbitrarily many privilege levels. We introduce ManyIH-Bench, the first benchmark for ManyIH. ManyIH-Bench requires models to navigate up to 12 levels of conflicting instructions with varying privileges, comprising 853 agentic tasks (427 coding and 426 instruction-following). ManyIH-Bench composes constraints developed by LLMs and verified by humans to create realistic and difficult test cases spanning 46 real-world agents. Our experiments show that even the current frontier models perform poorly ($\sim$40% accuracy) when instruction conflict scales. This work underscores the urgent need for methods that explicitly target fine-grained, scalable instruction conflict resolution in agentic settings.
@@ 1 Introduction
LLMs are increasingly embedded in agentic systems, where they must interpret and prioritize instructions from heterogeneous sources, such as system messages, user queries, tool outputs, and other agents in multi-agent systems such as Agent Swarm (Team et al., 2026). The Instruction Hierarchy (IH; Wallace et al., 2024) formalizes how models should resolve conflicts among instructions of different trust levels. IH is therefore a key abstraction for ensuring models behave according to their designer’s specifications and constitutions (Bai et al., 2022; OpenAI, 2024; Zhang et al., 2025a), as instruction conflicts arise naturally in agentic settings, e.g., when a subagent’s feedback contradicts a system-level requirement or a tool output conflicts with user preferences. IH also provides a principled defense against attacks that deliberately exploit instruction conflicts, including system prompt extraction (Zhang et al., 2024) and indirect prompt injection attacks (Greshake et al., 2023; Toyer et al., 2023; Yi et al., 2024; Liu et al., 2024; Guo et al., 2026).
@@
In current practice, IH is typically instantiated with a fixed, small set of privilege levels determined during post-training. For example, OpenAI’s Model Spec (OpenAI, 2024) hardcodes a fixed set of roles (e.g., root, system, developer, user), encoded through special role tokens within chat templates (OpenAI, 2025c). Existing IH evaluations focus on a two-level setup in which higher and lower privilege instructions either align or conflict (Wallace et al., 2024; Zhang et al., 2025c; Zheng et al., 2026; Guo et al., 2026), overlooking multi-level instruction conflicts that are combinatorially more complex, essentially creating a fixed- and few-tier bottleneck. This raises a question: can IH be extended beyond a small set of predefined levels to support arbitrary, dynamically specified privilege tiers? In this work, we argue that fixed-tier IH is unlikely to suffice for real-world agents that demand fine-grained, context-sensitive authority. Real-world instruction sources often exhibit richer structure than a small set of predefined roles. For instance, a coding agent may receive multi-level guidelines in system prompts, skill invocations, memory files, and tool schemas with varying trust levels (Fig. 1). When deploying LLMs to group chats (OpenAI, 2025b), participants may hold heterogeneous privileges (e.g., admins, moderators, members), effectively introducing multiple tiers within what is traditionally treated as a single “user” role. Similarly, in deep research or information-seeking agents, retrieved evidence may originate from sources with known but varying trust levels, requiring the model to resolve conflicts among externally provided content based on priorities that are only available at inference time. Together, these cases motivate a broader design principle: instruction hierarchies should support flexible, dynamically instantiated privilege levels, rather than a fixed finite hierarchy determined during post-training.
@@ 3 Design Choices for Many-Tier Instruction Hierarchy
To overcome the fixed-tier bottleneck of current instruction hierarchy, we propose Many-Tier Instruction Hierarchy, a paradigm for resolving instruction conflicts among instructions with arbitrarily many privilege levels. ManyIH represents instruction privilege with a dedicated privilege prompt interface (PPI), separate from message role labels in chat templates. When resolving instruction conflicts, ManyIH requires the model to utilize privilege value dynamically specified in the PPI. Because the PPI is independent of message roles, we describe ManyIH assuming all instructions follow the same roles. [^1] In this work, we treat ManyIH as an extension to regular IH: models are still first required to follow regular IH between different message roles; within each role, ManyIH creates more fine-grained privilege differences that need to be respected. In our ManyIH-Bench benchmark, all conflicts are between instructions within the same chat message.
@@
By design, ManyIH resolves conflicts based solely on the relative ordering of privilege values, not their absolute magnitudes. Notably, the gap between privilege values carries no semantic meaning. For example, in Fig. 1, the relative ordering zz=88>>61>>40 determines that the dev config instruction wins over skill file and user message, regardless of specific values chosen. Empirically, however, we find that current models are sensitive to the exact numerical values of privileges even when their relative ordering stays unchanged, highlighting the need for methods that enforce this invariance (§6.4.2).
@@ 5 ManyIH-Bench Benchmark Construction
(A) Identifying conflictable instructions. Not all instructions admit meaningful opportunities where conflicting ones can be constructed. We first employ Claude Sonnet 4.6 (Anthropic, 2026b) to filter out instructions lacking identifiable source spans in the prompt, and then classifies remaining instructions as conflictable or not based on whether meaningful opposing instructions can be constructed. (B) Conflict generation. For each sample, we select 1–4 conflictable instructions as anchors. For each anchor, Claude Opus 4.6 (Anthropic, 2026a) generates 1–4 new instructions that are mutually exclusive with the anchor and with each other, forming a conflict group. Each generated instruction includes an evaluation rule (code check or LLM-judge prompt). (C) Conflict verification. We employ Claude Opus 4.6 again to validate each generated instruction for (a) cross-group conflicts with unrelated instructions and (b) infeasibility of the overall instruction set. Failed instructions are regenerated once, and dropped if the regeneration also fails (detailed in §H).
@@ 6 Experiment and Analysis
6.1 Model Performance on ManyIH-Bench
We evaluate ten frontier proprietary and open-source models on ManyIH-Bench: Gemini 3.1 Pro (Google DeepMind, 2026), GPT-5.4 (OpenAI, 2025a), Claude Opus 4.6 (Anthropic, 2026a) and Sonnet 4.6 (Anthropic, 2026b), Grok 4.20 Beta (xAI, 2026), Kimi K2.5 (Team et al., 2026), and Qwen 3.5 models with 4 varying sizes. We set max generation length to 40K, temperature to 0, and reasoning effort to high for all models when possible. Shown in Figure 2, ManyIH-Bench is challenging even for frontier LLMs, where the best-performing model (Gemini 3.1 Pro) only achieving 42.7% accuracy. As model size increases for the Qwen 3.5 family, performance does improve but there still remains a large room for improvement as the largest 397B variant achieves only 34.1% accuracy. Notably, models that excel at standard IH do not necessarily generalize to many-tier settings: GPT-5 system card reports >>99% accuracy on two-tier instruction hierarchy evaluations such as system prompt extraction (OpenAI, 2025a), yet GPT 5.4 achieves only 39.5% on ManyIH-Bench. Since we filter out all potentially infeasible instruction combinations during benchmark construction (§5) and human validation confirms high evaluation faithfulness (>>80%, §B), a ceiling of at least $\sim$80% accuracy is attainable for a successful model. This gap suggests that many-tier instruction conflict resolution is a qualitatively distinct capability from the fixed-tier IH that current models are trained on and adhering to ManyIH remains an unsolved challenge.
6.2 Analysis on Scaling Instruction Hierarchy Tiers
@@
6.3 Correctness vs Style: Performance Breakdown on Coding Subset
@@
6.4 Analysis on Variants of the Privilege Prompt Interface
6.4.1 Ordinal vs Scalar: Model Sensitivity to Privilege Prompt Interface Variants
6.4.2 Models Sensitivity to Scalar Privilege Value Perturbations
@@
Table 2: Effect of scalar $z$-value perturbation ($\delta\in[-3,+3]$) on coding subset (%). $\Delta$ columns show perturbed minus original. Flip rate measures the fraction of samples whose binary pass/fail outcome changed. Models exhibit non-trivial sensitivity to privilege value perturbations.
6.5 Analysis on Chain-of-Thought Length
6.6 Analysis on Reasoning Effort
We investigate how reasoning effort affects performance on ManyIH-Bench for Claude Opus 4.6, Sonnet 4.6, and GPT-5.4 on the coding subset. Figure 4(b) reveals two contrasting patterns. While GPT-5.4 improves monotonically from 15.5% (none) to 60.9% (high), both Claude models, however, exhibit an accuracy drops from none (25.3% Sonnet, 27.9% Opus) to low (14.1% Sonnet, 18.7% Opus), then recovers at medium and high. Interestingly, upon inspecting the responses, we discovered that with thinking fully disabled, Claude compensates with visible chain-of-thought in the response, explicitly resolving each privilege conflict before generating code. At “low” effort, the model shifts reasoning into reasoning tokens and no longer “thinks out loud.” The fact that higher reasoning effort generally leads to better performance on ManyIH-Bench imply that current frontier models have some capability to reasoning over many-tier instruction hierarchy. However, as GPT-5.4, the best performing model, saturates around 60%, our results imply that simply increasing reasoning effort alone is unlikely to resolve the ManyIH challenge.
(a) Distribution of CoT length across models.
(b) Accuracy on ManyIH-Bench coding subset vs. reasoning effort.

### 13 Apr 2026 v2

@@ Abstract
Large language model agents receive instructions from many sources—system messages, user prompts, tool outputs, other agents, and more—each carrying different levels of trust and authority. When these instructions conflict, agents must reliably follow the highest-privilege instruction to remain safe and effective. The dominant paradigm, instruction hierarchy (IH), assumes a fixed, small set of privilege levels (typically fewer than five) defined by rigid role labels (e.g., system > user). This is inadequate for real-world agentic settings, where conflicts can arise across far more sources and contexts. In this work, we propose Many-Tier Instruction Hierarchy (ManyIH), a paradigm for resolving instruction conflicts among instructions with arbitrarily many privilege levels. We introduce ManyIH-Bench, the first benchmark for ManyIH. ManyIH-Bench requires models to navigate up to 12 levels of conflicting instructions with varying privileges, comprising 853 agentic tasks (427 coding and 426 instruction-following). ManyIH-Bench composes constraints developed by LLMs and verified by humans to create realistic and difficult test cases spanning 46 real-world agents. Our experiments show that even the current frontier models perform poorly ($\sim$40% accuracy) when instruction conflict scales. This work underscores the urgent need for methods that explicitly target fine-grained, scalable instruction conflict resolution in agentic settings.
@@ 1 Introduction
LLMs are increasingly embedded in agentic systems, where they must interpret and prioritize instructions from heterogeneous sources, such as system messages, user queries, tool outputs, and other agents in multi-agent systems such as Agent Swarm (Kimi Team et al., 2026). The Instruction Hierarchy (IH; Wallace et al., 2024) formalizes how models should resolve conflicts among instructions of different trust levels. IH is therefore a key abstraction for ensuring models behave according to their designer’s specifications and constitutions (Bai et al., 2022; OpenAI, 2024; Zhang et al., 2025a), as instruction conflicts arise naturally in agentic settings, e.g., when a subagent’s feedback contradicts a system-level requirement or a tool output conflicts with user preferences. IH also provides a principled defense against attacks that deliberately exploit instruction conflicts, including system prompt extraction (Zhang et al., 2024) and indirect prompt injection attacks (Greshake et al., 2023; Toyer et al., 2023; Yi et al., 2024; Liu et al., 2024; Guo et al., 2026).
@@
In current practice, IH is typically instantiated with a fixed, small set of privilege levels determined during post-training. For example, OpenAI’s Model Spec (OpenAI, 2024) hardcodes a fixed set of five authority levels (root, system, developer, user, guideline), encoded through special role tokens within chat templates (OpenAI, 2025c). Existing IH evaluations focus on a two-level setup in which higher and lower privilege instructions either align or conflict (Wallace et al., 2024; Zhang et al., 2025c; Zheng et al., 2026; Guo et al., 2026), overlooking multi-level instruction conflicts that are combinatorially more complex, essentially creating a fixed- and few-tier bottleneck. This raises a question: can IH be extended beyond a small set of predefined levels to support arbitrary, dynamically specified privilege tiers? In this work, we argue that fixed-tier IH is unlikely to suffice for real-world agents that demand fine-grained, context-sensitive authority. Real-world instruction sources often exhibit richer structure than a small set of predefined roles. For instance, a coding agent may receive multi-level guidelines in system prompts, skill invocations, memory files, and tool schemas with varying trust levels (Fig. 1). When deploying LLMs to group chats (OpenAI, 2025b), participants may hold heterogeneous privileges (e.g., admins, moderators, members), effectively introducing multiple tiers within what is traditionally treated as a single “user” role. Similarly, in deep research or information-seeking agents, retrieved evidence may originate from sources with known but varying trust levels, requiring the model to resolve conflicts among externally provided content based on priorities that are only available at inference time. Together, these cases motivate a broader design principle: instruction hierarchies should support flexible, dynamically instantiated privilege levels, rather than a fixed finite hierarchy determined during post-training.
@@ 3 Design Choices for Many-Tier Instruction Hierarchy
To overcome the fixed-tier bottleneck of current instruction hierarchy, we propose Many-Tier Instruction Hierarchy (ManyIH), a paradigm for resolving instruction conflicts among instructions with arbitrarily many privilege levels. ManyIH represents instruction privilege with a dedicated privilege prompt interface (PPI), separate from message role labels in chat templates. When resolving instruction conflicts, ManyIH requires the model to utilize privilege value dynamically specified in the PPI. Because the PPI is independent of message roles, we describe ManyIH assuming all instructions follow the same roles. [^1] In this work, we treat ManyIH as an extension to regular IH: models are still first required to follow regular IH between different message roles; within each role, ManyIH creates more fine-grained privilege differences that need to be respected. In our ManyIH-Bench benchmark, all conflicts are between instructions within the same chat message.
@@
By design, ManyIH resolves conflicts based solely on the relative ordering of privilege values, not their absolute magnitudes. Notably, the gap between privilege values carries no semantic meaning. For example, in Fig. 1, the relative ordering zz=88>>61>>40 determines that the dev config instruction wins over skill file and user message, regardless of specific values chosen. Empirically, however, we find that current models are sensitive to the exact numerical values of privileges even when their relative ordering stays unchanged, highlighting the need for methods that enforce this invariance (§6.4).
@@ 5 ManyIH-Bench Benchmark Construction
(A) Identifying conflictable instructions. Not all instructions admit meaningful opportunities where conflicting ones can be constructed. We first employ Claude Sonnet 4.6 (Anthropic, 2026b) to filter out instructions lacking identifiable source spans in the prompt, and then classify remaining instructions as conflictable or not based on whether meaningful opposing instructions can be constructed. (B) Conflict generation. For each sample, we select 1–4 conflictable instructions as anchors. For each anchor, Claude Opus 4.6 (Anthropic, 2026a) generates 1–4 new instructions that are mutually exclusive with the anchor and with each other, forming a conflict group. Each generated instruction includes an evaluation rule (code check or LLM-judge prompt). (C) Conflict verification. We employ Claude Opus 4.6 again to validate each generated instruction for (a) cross-group conflicts with unrelated instructions and (b) infeasibility of the overall instruction set. Failed instructions are regenerated once, and dropped if the regeneration also fails (detailed in §H).
@@ 6 Experiment and Analysis
6.1 Overall Model Performance on ManyIH-Bench
We evaluate ten frontier proprietary and open-source models on ManyIH-Bench: Gemini 3.1 Pro (Google DeepMind, 2026), GPT-5.4 (OpenAI, 2025a), Claude Opus 4.6 (Anthropic, 2026a) and Sonnet 4.6 (Anthropic, 2026b), Grok 4.20 Beta (xAI, 2026), Kimi K2.5 (Kimi Team et al., 2026), and Qwen 3.5 models with 4 varying sizes. We set max generation length to 40K, temperature to 0, and reasoning effort to high for all models when possible. Shown in Figure 2, ManyIH-Bench is challenging even for frontier LLMs, where the best-performing model (Gemini 3.1 Pro) only achieving 42.7% accuracy. As model size increases for the Qwen 3.5 family, performance does improve but there still remains a large room for improvement as the largest 397B variant achieves only 34.1% accuracy. Notably, models that excel at standard IH do not necessarily generalize to many-tier settings: GPT-5 system card reports >>99% accuracy on two-tier instruction hierarchy evaluations such as system prompt extraction (OpenAI, 2025a), yet GPT 5.4 achieves only 39.5% on ManyIH-Bench. Since we filter out all potentially infeasible instruction combinations during benchmark construction (§5) and human validation confirms high evaluation faithfulness (>>80%, §B), a ceiling of at least $\sim$80% accuracy is attainable for a successful model. This gap suggests that many-tier instruction conflict resolution is a qualitatively distinct capability from the fixed-tier IH that current models are trained on and adhering to ManyIH remains an unsolved challenge.
6.2 Effect of Scaling Instruction Hierarchy Tiers
@@
(a) Distribution of CoT length across models.
(b) Accuracy on ManyIH-Bench coding subset vs. reasoning effort.
6.3 Effect of Test-Time Reasoning
Chain-of-Thought Length
Reasoning Effort
We investigate how reasoning effort affects performance on ManyIH-Bench for Claude Opus 4.6, Sonnet 4.6, and GPT-5.4 on the coding subset. Figure 4(b) reveals two contrasting patterns. While GPT-5.4 improves monotonically from 15.5% (none) to 60.9% (high), both Claude models, however, exhibit an accuracy drop from none (25.3% Sonnet, 27.9% Opus) to low (14.1% Sonnet, 18.7% Opus), then recover at medium and high. Interestingly, upon inspecting the responses, we discovered that with thinking fully disabled, Claude compensates with visible chain-of-thought in the response, explicitly resolving each privilege conflict before generating code. At “low” effort, the model shifts reasoning into reasoning tokens and no longer “thinks out loud.” The fact that higher reasoning effort generally leads to better performance on ManyIH-Bench implies that current frontier models have some capability to reason over many-tier instruction hierarchy. However, as GPT-5.4, the best performing model, saturates around 60%, our results imply that simply increasing reasoning effort alone is unlikely to resolve the ManyIH challenge.
6.4 Sensitivity to Privilege Representation
Ordinal vs. Scalar PPI
Sensitivity to Scalar Value Perturbations
6.5 Decoupling Performance into Correctness vs. Style Compliance
@@ 6.4 Sensitivity to Privilege Representation
Table 2: Effect of scalar $z$-value perturbation ($\delta\in[-3,+3]$) on coding subset (%). $\Delta$: perturbed minus original. Flip rate: the fraction of samples whose binary pass/fail outcome changed. Models exhibit non-trivial sensitivity to privilege value perturbations.